OpenStudy (chycora):

Estimate the domain of the following graph. A. x > 0 B. 0 x 7 C. -3 x 0 D. x is all real numbers.

OpenStudy (chycora):

is the answer B

OpenStudy (chycora):

and b and c are B- \[0 \le x \le 7\] C- \[-3 \le x \le 0\]

OpenStudy (chycora):

@mathstudent55

OpenStudy (peachpi):

no, the range (y-values) are from 0 to 7. The domain is the x-values. Can you tell what the smallest and largest x values are?

OpenStudy (animaluver15):

b is correct

OpenStudy (peachpi):

b is not correct

OpenStudy (chycora):

why not

OpenStudy (peachpi):

I explained above

OpenStudy (chycora):

so would it be either A or D

OpenStudy (chycora):

@563blackghost

OpenStudy (chycora):

@jabez177

OpenStudy (sstradic):

A

OpenStudy (chycora):

ok thanks

OpenStudy (chycora):

it was C